OpenStudy (anonymous):

For the function f(x) = 2 - 3x, find f(0).

geerky42 (geerky42):

\[f(\mathbf{\color{red}x}) = 2-3\mathbf{\color{red}x} \\~\\f(\mathbf{\color{red}0}) = 2-3(\mathbf{\color{red}0})=\cdots~?\]

OpenStudy (anonymous):

-1 @geerky42

OpenStudy (anonymous):

?

geerky42 (geerky42):

What is 3*0 ?